OpenStudy (anonymous):

For each function,find f(-2),f(-1/2),f(3) and f(7) f(x)=-3x+1.8

OpenStudy (anonymous):

f(-2)=3(-2)+1.8 f(-1/2)=3(-1/2)+1.8 f(3)=3(3)+1.8 f(7)=3(7)+1.8

OpenStudy (anonymous):

f(-2) = 7.8 f(-1/2) = 3.3 f(3) = -7.2 f(7) = 19.2

OpenStudy (precal):

f(7) =-19.2

OpenStudy (anonymous):

ya, it is negati9ve, my bad. :-)